# Does it matter? A more careful treatment of density fluctuations in 21-cm simulations

### v1

@@ Abstract
The cosmological 21-cm signal is sourced from hyperfine transitions in neutral hydrogen atoms. Yet, although hydrogen atoms belong to the sub-class of baryonic matter, it is customary to simplify their treatment and analyze the 21-cm signal as if all the matter in the Universe was in the form of collisionless cold dark matter (CDM). This is usually done by evolving the density field via a scale-independent growth factor (SIGF). In this work, we separate the baryons from CDM and evolve the two species with a proper scale-dependent growth factor (SDGF). By incorporating the SDGF in the 21cmFirstCLASS code, we demonstrate the effect that baryons and CDM have on the 21-cm signal at the linear dark ages epoch and the subsequent non-linear epochs of cosmic dawn and reionization. Our analysis shows that the baryonic nature of hydrogen cannot be ignored during the dark ages, and that non-linear effects in density-field evolution must be accounted for after stars have formed. Furthermore, we discuss how the 21-cm signal is modified at lower redshifts, where ground-based 21-cm interferometers are mostly sensitive, due to the choice of working with either the “linear” or “non-linear” matter over-density (that is, the over-density as computed from linear perturbation theory, and non-linear perturbation theory, respectively) in the extended Press-Schechter formalism. Our code is publicly available at https://github.com/jordanflitter/21cmFirstCLASS.
@@ I Introduction
Between recombination and reionization, the Universe was filled with neutral hydrogen atoms. Hyperfine energy de-excitations in these atoms, following interactions with other particles, led to the emission of 21-cm wavelength photons. This is the 21-cm signal in a nutshell Madau et al. (1997); Barkana and Loeb (2001); Loeb and Zaldarriaga (2004); Bharadwaj and Ali (2005); Furlanetto et al. (2006); Pritchard and Loeb (2012); Bera et al. (2023); Shaw et al. (2023). This exciting signal has the potential to expand our knowledge on the formation mechanism of the first galaxies Barkana and Loeb (2005a); Furlanetto (2006); Park et al. (2019); Muñoz et al. (2020); Park et al. (2020); Greig et al. (2020a); Thyagarajan et al. (2020); Qin et al. (2020, 2021a); Reis et al. (2022a); Ma et al. (2023), the thermal history of the intergalactic medium (IGM) Chen and Miralda-Escude (2004); Mesinger et al. (2014); Ghara and Mellema (2020); Greig et al. (2021); Ghara et al. (2021); Maity and Choudhury (2022); Abdurashidova et al. (2022, 2023); Lazare et al. (2024a), the epoch of reionization Lidz et al. (2009); Wiersma et al. (2013); Jensen et al. (2013); Meerburg et al. (2013); Greig and Mesinger (2015, 2017); Greig et al. (2020b); Ghara et al. (2020); Watkinson et al. (2022); Rahimi et al. (2021); Qin et al. (2021b); Saxena et al. (2023); Giri et al. (2024); Ghara et al. (2024), physics beyond the standard model Tashiro and Sugiyama (2006); Schleicher et al. (2009); Shiraishi et al. (2014); Chluba et al. (2015); Villaescusa-Navarro et al. (2015); Oyama et al. (2016); Pal and Guha Sarkar (2016); Kovetz et al. (2019); Caputo et al. (2021); Natwariya and Nayak (2022); Kunze (2019); Minoda et al. (2022); Shmueli et al. (2023); Kunze (2023); Cruz et al. (2024a); Libanore et al. (2024); Bhaumik et al. (2024); Lee and Hotinli (2024); Dhuria and Teli (2024); Plombat et al. (2024); Adi et al. (2024), and possibly shed some light on the nature of dark matter Tashiro et al. (2014); Evoli et al. (2014); Muñoz et al. (2015a); Lopez-Honorez et al. (2016); D’Amico et al. (2018); Cheung et al. (2019); Mitridate and Podo (2018); Lidz and Hui (2018); Nebrin et al. (2019); Fialkov et al. (2018); Barkana (2018); Clark et al. (2018); Kovetz et al. (2018); Berlin et al. (2018); Barkana et al. (2018); Slatyer and Wu (2018); Muñoz and Loeb (2018); Liu and Slatyer (2018); Muñoz et al. (2018); Creque-Sarbinowski et al. (2019); Liu et al. (2019); Hotinli et al. (2022); Jones et al. (2021); Halder et al. (2021); Sarkar et al. (2022); Barkana et al. (2023); Giri and Schneider (2022); Flitter and Kovetz (2022); Lazare et al. (2024b); Driskell et al. (2022); Dey et al. (2023); Short et al. (2022); Mondal et al. (2023); Facchinetti et al. (2024); Qin et al. (2024); Sun et al. (2023); Vanzan et al. (2024); Cang et al. (2023); Shao et al. (2023); Katz et al. (2024); Lopez-Honorez (2024).
@@
For these reasons, worldwide efforts are being made to measure the 21-cm signal by various collaborations and experiments. Some of these experiments are designed to measure the “global” (sky-averaged) 21-cm signal, like the Experiment to Detect the Global reionization Signature (EDGES) Monsalve et al. (2019), Shaped Antenna measurement of the background RAdio Spectrum (SARAS) Nambissan T. et al. (2021), Large-Aperture Experiment to Detect the Dark Ages (LEDA) Price et al. (2018), the Radio Experiment for the Analysis of Cosmic Hydrogen (REACH) de Lera Acedo et al. (2022) and Probing Radio Intensity at high-Z from Marion (PRizM) Philip et al. (2019). Other experiments, like the Hydrogen Epoch of Reionization Array (HERA) DeBoer et al. (2017), Low Frequency Array (LOFAR) van Haarlem et al. (2013), Murchison Widefield Array (MWA) Tingay et al. (2013), the Giant Metrewave Radio Telescope (GMRT) Swarup et al. (1991), the Precision Array for Probing the Epoch of Reionization (PAPER) Kolopanis et al. (2019) and Square Kilometre Array (SKA) Braun et al. (2015) have been dedicated to measuring the spatial fluctuations in the signal using ground based interferometers. All the aforementioned experiments are limited to detecting the signal at redshifts $z\lesssim 30$ (namely, high frequencies), as the Earth ionosphere distorts incoming low-frequency radiation that is sourced at very high redshifts. Several interesting proposals are considered nowadays to overcome this obstacle by deploying radio receivers on the far side of the Moon, or in orbit around it or the Earth Jester and Falcke (2009); Burns et al. (2021a); Furlanetto et al. (2019); Silk (2021); Burns et al. (2021b, a); Goel et al. (2022); Shi et al. (2022); Bertone et al. (2023); Fialkov et al. (2023).
@@
The growing interest in the 21-cm signal and its analysis motivated the development of a plethora of dedicated codes, designed to estimate the signal and its spatial fluctuations as a function of redshift. After recombination, prior to star formation, the fluctuations in the signal are tiny and linear perturbation theory can be applied in that regime. For example, CAMB Lewis and Challinor (2007) and PowerSpectrum21cm Muñoz et al. (2015b) exploit the linearity of the perturbations in this “dark ages” epoch. After cosmic dawn, once galaxies and stars have formed, the fluctuations in the signal are amplified considerably, and the validity of linear perturbation theory breaks. Different techniques have been adapted in the literature to simulate the signal in the non-linear regime of cosmic dawn and reionization. This includes full radiative transfer simulations (CodA Ocvirk et al. (2016, 2020); Lewis et al. (2022), 21SSD Semelin et al. (2017), THESAN Kannan et al. (2022)), one-dimensional radiative transfer methods (BEoRN Schaeffer et al. (2023), Bears Thomas et al. (2009), Grizzly Ghara et al. (2018)), hydrodynamic-radiative-transfer frameworks (Licorice Semelin et al. (2007); Meriot and Semelin (2024)), and ray-tracing algorithms (C2-ray Mellema et al. (2006); Hirling et al. (2023), CRASH Maselli et al. (2003)).
Figure 2: Comparison of the 21-cm power spectrum on different scales, under the same conditions as in Fig. 1. Here, we show with error bars the expected noise from HERA in its design sensitivity at low redshifts ($z<10$) assuming the moderate foreground scenario, as evaluated by the 21cmSense code Pober et al. (2013, 2014) (we used the same 21cmSense settings as detailed in Ref. Flitter and Kovetz (2024a)). We also show a separate error bar at $z=14$, corresponding to a sensitivity estimate for SKA Meriot and Semelin (2024).
@@
Semi-numerical codes, whose runtime is shorter by orders of magnitude, are also found in the literature. Among these, 21cmFAST Mesinger et al. (2011); Muñoz et al. (2022) is perhaps the most popular one, having also been used to analyse HERA Phase I data Abdurashidova et al. (2022, 2023), but see also simfast21 Santos et al. (2010), CIFOG Hutter (2018) and 21cmSPACE Visbal et al. (2012); Fialkov et al. (2013, 2014a, 2014b); Fialkov and Barkana (2014); Fialkov et al. (2015); Cohen et al. (2016); Fialkov et al. (2017); Fialkov and Barkana (2019); Reis et al. (2020, 2021, 2022b); Magg et al. (2022); Gessey-Jones et al. (2022); Sikder et al. (2023); Gessey-Jones et al. (2023); Pochinda et al. (2024), though the latter is not publicly available as of yet. Recently, using output from 21cmFAST, and assuming linear evolution of the density field, a fully analytic code named Zeus21 Muñoz (2023); Cruz et al. (2024b) was calibrated at the cosmic dawn era, allowing the code to compute the 21-cm power spectrum in mere seconds. [^1] ARES Mirocha (2014); Mirocha et al. (2017) is another speedy code, but unlike Zeus21 and HMreio, it only computes the global 21-cm signal. Another analytic code, named HMreio Schneider et al. (2023), also calibrated to 21cmFAST, focuses on the epoch of reionization. All these codes use a single quantity $\delta$ to denote over-density, and thus implicitly assume $\delta_{b}=\delta_{c}=\delta_{m}$, namely that baryons (“$b$”) and cold dark matter (CDM, “$c$”) behave alike (“$m$” stands for all of matter).
@@
In this paper, we relax this assumption and explore the cosmological implications of distinguishing $\delta_{b}$ from $\delta_{c}$ in the context of the 21-cm signal. To do so, we rely on the newly developed tool 21cmFirstCLASS Flitter and Kovetz (2024a, b). This code interfaces CLASS Blas et al. (2011) and 21cmFAST, allowing the simulation to begin from recombination with initial conditions that are consistent with the simulated cosmology. We find that at the dark ages the approximation $\delta_{b}\approx\delta_{c}$ completely fails, and $\delta_{b}$ must be evolved in a scale-dependent manner in that epoch. We also demonstrate, using scale-dependent Lagrangian perturbation theory (LPT), the need to go beyond linear perturbation theory for the baryon over-density field at low redshifts, after cosmic dawn was set. Finally, we present how the signal is modified when considering either the “linear” or “non-linear” matter over-density $\delta_{m}$ in the context of the extended Press-Schechter formalism, and we discuss the role of the relative velocity between baryons and dark matter in determining the 21-cm signal. Our main results are shown in Figs. 1-2.
In Fig. 2, the $z\lesssim 10$ error bars indicate the expected noise level from HERA in its design sensitivity under the conservative assumption of a moderate foreground scenario, evaluated with 21cmSense Pober et al. (2013, 2014), while the $z=14$ error bar illustrates the sensitivity expected for SKA, estimated by Ref. Meriot and Semelin (2024). This figure demonstrates that the effects discussed in this work should matter in future analyses of 21-cm signal measurements.
@@
This paper is organized as follows. In Sec. II, we examine the impact of using the baryon density in the computation of the 21-cm signal during the dark ages. We then move to analyzing the role of the baryons and CDM during cosmic dawn and reionization in Sec. III. We discuss in Sec. IV our results from the previous section and write our conclusions in Sec. V. In addition, this paper contains several appendices. In Appendix A, we study the sensitivity of the 21-cm signal to the precise evolution of CDM and the total matter density fields. In Appendix B, we study the Boltzmann equation for the spin temperature. In Appendix C, we verify the output of 21cmFirstCLASS, by comparing the temperature fluctuations at the dark ages with the analytical prediction from linear perturbation theory. Finally, we derive the equations for second-order scale-dependent LPT in Appendix D.
@@
II Dark ages
II.1 Matter density fields at the dark ages
According to the standard cosmological model, two kinds of matter controlled the evolution of the Universe at early times, before the first atoms were formed. The first kind are baryons, which include hydrogen and helium nuclei and electrons and protons that interact strongly with photons due to their electric charge. The second kind is collisionless CDM. While the baryons were subject to radiation pressure on small scales, CDM was able to freely cluster under the influence of self-gravitational forces. As the Universe expanded, it eventually reached a low enough temperature such that atoms could have survived immediate ionization for the first time. This promptly led to a dramatic increase in the photon mean free path, thus releasing the freely propagating cosmic microwave background (CMB) radiation. This key moment in the history of the Universe, dated at redshift $z\approx 1100$, is known as “recombination”.

### v2

@@ Abstract
The cosmological 21-cm signal is sourced from hyperfine transitions in neutral hydrogen atoms. Yet, although the abundance of hydrogen atoms follows the baryon density field, semi-numerical codes that simulate the 21-cm signal simplify their treatment as if all the matter in the Universe was in the form of collisionless cold dark matter (CDM). This is usually done by evolving the density field via a scale-independent growth factor (SIGF). In this work, we separate the baryons from CDM and evolve the two species with a proper scale-dependent growth factor (SDGF). By incorporating the SDGF in the 21cmFirstCLASS code, we demonstrate the effect that baryons and CDM have on the 21-cm signal at the linear dark ages epoch and the subsequent non-linear epochs of cosmic dawn and reionization. Our analysis shows that the baryonic nature of hydrogen cannot be ignored during the dark ages, and that non-linear effects in density-field evolution must be accounted for after stars have formed. Furthermore, we discuss how the 21-cm signal is modified at lower redshifts, where ground-based 21-cm interferometers are mostly sensitive, due to the choice of working with either the “linear” or “non-linear” matter over-density (that is, the over-density as computed from linear perturbation theory, and non-linear perturbation theory, respectively) in the extended Press-Schechter formalism. Our code is publicly available at https://github.com/jordanflitter/21cmFirstCLASS.
@@ I Introduction
Between recombination and reionization, the Universe was filled with neutral hydrogen atoms. Hyperfine energy de-excitations in these atoms, following interactions with other particles, led to the emission of 21-cm wavelength photons. This is the 21-cm signal in a nutshell Madau et al. (1997); Barkana and Loeb (2001); Loeb and Zaldarriaga (2004); Bharadwaj and Ali (2005); Furlanetto et al. (2006); Pritchard and Loeb (2012); Bera et al. (2023); Shaw et al. (2023). This exciting signal has the potential to expand our knowledge on the formation mechanism of the first galaxies Barkana and Loeb (2005a); Furlanetto (2006); Park et al. (2019); Muñoz et al. (2020); Park et al. (2020); Greig et al. (2020a); Thyagarajan et al. (2020); Qin et al. (2020, 2021a); Reis et al. (2022a); Ma et al. (2023), the thermal history of the intergalactic medium (IGM) Chen and Miralda-Escude (2004); Mesinger et al. (2014); Ghara and Mellema (2020); Greig et al. (2021); Mittal and Kulkarni (2021); Ghara et al. (2021); Maity and Choudhury (2022); Abdurashidova et al. (2022, 2023); Lazare et al. (2024a), the epoch of reionization Lidz et al. (2009); Wiersma et al. (2013); Jensen et al. (2013); Meerburg et al. (2013); Greig and Mesinger (2015, 2017); Greig et al. (2020b); Ghara et al. (2020); Watkinson et al. (2022); Rahimi et al. (2021); Qin et al. (2021b); Saxena et al. (2023); Giri et al. (2024); Ghara et al. (2024), physics beyond the standard model Tashiro and Sugiyama (2006); Schleicher et al. (2009); Shiraishi et al. (2014); Chluba et al. (2015); Villaescusa-Navarro et al. (2015); Oyama et al. (2016); Pal and Guha Sarkar (2016); Kovetz et al. (2019); Caputo et al. (2021); Natwariya and Nayak (2022); Kunze (2019); Minoda et al. (2022); Shmueli et al. (2023); Kunze (2023); Cruz et al. (2024a); Libanore et al. (2024); Bhaumik et al. (2025); Lee and Hotinli (2024); Dhuria and Teli (2024); Plombat et al. (2024); Adi et al. (2024); Libanore et al. (2025), and possibly shed some light on the nature of dark matter Tashiro et al. (2014); Evoli et al. (2014); Muñoz et al. (2015a); Lopez-Honorez et al. (2016); D’Amico et al. (2018); Cheung et al. (2019); Mitridate and Podo (2018); Lidz and Hui (2018); Nebrin et al. (2019); Fialkov et al. (2018); Barkana (2018); Clark et al. (2018); Kovetz et al. (2018); Berlin et al. (2018); Barkana et al. (2018); Slatyer and Wu (2018); Muñoz and Loeb (2018); Liu and Slatyer (2018); Muñoz et al. (2018); Creque-Sarbinowski et al. (2019); Liu et al. (2019); Hotinli et al. (2022); Jones et al. (2021); Halder et al. (2021); Mittal et al. (2022); Sarkar et al. (2022); Barkana et al. (2023); Giri and Schneider (2022); Flitter and Kovetz (2022); Lazare et al. (2024b); Driskell et al. (2022); Dey et al. (2023); Short et al. (2022); Mondal et al. (2023); Facchinetti et al. (2024); Qin et al. (2024); Sun et al. (2023); Vanzan et al. (2024); Cang et al. (2023); Shao et al. (2023); Katz et al. (2024); Lopez-Honorez (2024).
@@
For these reasons, worldwide efforts are being made to measure the 21-cm signal by various collaborations and experiments. Some of these experiments are designed to measure the “global” (sky-averaged) 21-cm signal, like the Experiment to Detect the Global reionization Signature (EDGES) Monsalve et al. (2019), Shaped Antenna measurement of the background RAdio Spectrum (SARAS) Nambissan T. et al. (2021), Large-Aperture Experiment to Detect the Dark Ages (LEDA) Price et al. (2018), the Radio Experiment for the Analysis of Cosmic Hydrogen (REACH) de Lera Acedo et al. (2022) and Probing Radio Intensity at high-Z from Marion (PRizM) Philip et al. (2019). Other experiments, like the Hydrogen Epoch of Reionization Array (HERA) DeBoer et al. (2017), Low Frequency Array (LOFAR) van Haarlem et al. (2013), Murchison Widefield Array (MWA) Tingay et al. (2013), the Giant Metrewave Radio Telescope (GMRT) Swarup et al. (1991), the Precision Array for Probing the Epoch of Reionization (PAPER) Kolopanis et al. (2019) and Square Kilometre Array (SKA) Braun et al. (2015) have been dedicated to measuring the spatial fluctuations in the signal using ground based interferometers. All the aforementioned experiments are limited to detecting the signal at redshifts $z\lesssim 30$ (namely, high frequencies), as the Earth ionosphere distorts incoming low-frequency radiation that is sourced at very high redshifts. Several interesting proposals are considered nowadays to overcome this obstacle by deploying radio receivers on the far side of the Moon, or in orbit around it or the Earth Jester and Falcke (2009); Burns et al. (2021a); Furlanetto et al. (2019); Silk (2021); Burns et al. (2021b, a); Goel et al. (2022); Shi et al. (2022); Bertone et al. (2023); Fialkov et al. (2023); Mondal and Barkana (2023).
@@
The growing interest in the 21-cm signal and its analysis motivated the development of a plethora of dedicated codes, designed to estimate the signal and its spatial fluctuations as a function of redshift. After recombination, prior to star formation, the fluctuations in the signal are tiny and linear perturbation theory can be applied in that regime. For example, CAMB Lewis and Challinor (2007) and PowerSpectrum21cm Muñoz et al. (2015b) exploit the linearity of the perturbations in this “dark ages” epoch, see also Ref. Naoz and Barkana (2005). After cosmic dawn, once galaxies and stars have formed, the fluctuations in the signal are amplified considerably, and the validity of linear perturbation theory breaks. Different techniques have been adapted in the literature to simulate the signal in the non-linear regime of cosmic dawn and reionization. This includes full radiative transfer simulations (CodA Ocvirk et al. (2016, 2020); Lewis et al. (2022), 21SSD Semelin et al. (2017), THESAN Kannan et al. (2022)), one-dimensional radiative transfer methods (BEoRN Schaeffer et al. (2023), Bears Thomas et al. (2009), Grizzly Ghara et al. (2018)), hydrodynamic-radiative-transfer frameworks (Licorice Semelin et al. (2007); Meriot and Semelin (2024)), and ray-tracing algorithms (C2-ray Mellema et al. (2006); Hirling et al. (2023), CRASH Maselli et al. (2003)).
@@
Semi-numerical codes, whose runtime is shorter by orders of magnitude, are also found in the literature. Among these, 21cmFAST Mesinger et al. (2011); Muñoz et al. (2022) is perhaps the most popular one, but see also simfast21 Santos et al. (2010), CIFOG Hutter (2018) and 21cmSPACE Visbal et al. (2012); Fialkov et al. (2013, 2014a, 2014b); Fialkov and Barkana (2014); Fialkov et al. (2015); Cohen et al. (2016); Fialkov et al. (2017); Fialkov and Barkana (2019); Reis et al. (2020, 2021, 2022b); Magg et al. (2022); Gessey-Jones et al. (2022); Sikder et al. (2023); Gessey-Jones et al. (2023); Pochinda et al. (2024), though the latter is not publicly available as of yet. Recently, using output from 21cmFAST, and assuming linear evolution of the density field, a fully analytic code named Zeus21 Muñoz (2023); Cruz et al. (2024b) was calibrated at the cosmic dawn era, allowing the code to compute the 21-cm power spectrum in mere seconds. [^1] ARES Mirocha (2014); Mirocha et al. (2017) and ECHO21 Mittal (2025) are also speedy codes, but unlike Zeus21 and HMreio, they only compute the global 21-cm signal. Another analytic code, named HMreio Schneider et al. (2023), also calibrated to 21cmFAST, focuses on the epoch of reionization. All these codes use a single quantity $\delta$ to denote over-density, and thus implicitly assume $\delta_{b}=\delta_{c}=\delta_{m}$, namely that baryons (“$b$”) and cold dark matter (CDM, “$c$”) behave alike (“$m$” stands for all of matter).
@@
In this paper, we relax this assumption and explore the cosmological implications of distinguishing $\delta_{b}$ from $\delta_{c}$ in the context of the 21-cm signal. To do so, we rely on the newly developed tool 21cmFirstCLASS Flitter and Kovetz (2024a, b). We find that at the dark ages the approximation $\delta_{b}\approx\delta_{c}$ fails when high order statistics (e.g. the power spectrum) are considered, and $\delta_{b}$ must be evolved in a scale-dependent manner in that epoch. We also demonstrate, using scale-dependent Lagrangian perturbation theory (LPT), the need to go beyond linear perturbation theory for the baryon over-density field at low redshifts, after cosmic dawn was set. Finally, we present how the signal is modified when considering either the “linear” or “non-linear” matter over-density $\delta_{m}$ in the context of the extended Press-Schechter formalism. The modifications made in 21cmFirstCLASS for this work now make it the only public code in the literature that can consistently simulate the fluctuations in the 21-cm signal at all cosmological epochs, from dark ages until reionization, while simultaneously studying the cosmic microwave background (CMB) anisotropies.
@@
This paper is organized as follows. Sec. II provides the theoretical background for the 21-cm signal and the matter fields that govern its shape. In Sec. III we give a short description of the main tool used in our analysis, 21cmFirstCLASS, and the modifications made for this work. We show our results in Sec. IV, discuss their significance in Sec. V, and conclude in Sec. VI. In addition, this paper contains several appendices. In Appendix A, we study the Boltzmann equation for the spin temperature. In Appendix B, we study the sensitivity of the 21-cm signal to the precise evolution of CDM and the total matter density fields. We derive the equations for second-order scale-dependent LPT in Appendix C. Finally, in Appendix D, we verify the output of 21cmFirstCLASS, by comparing the temperature fluctuations at the dark ages with the analytical prediction from linear perturbation theory.
@@
II Theory
In this section, we summarize the main theoretical components needed to consistently model the evolution of the 21-cm signal during the dark ages and cosmic dawn. All of them, have been implemented in the public code 21cmFirstCLASS, as described in Sec. III.
II.1 The 21-cm signal
[…]
The quantity $T_{s}$ that appears in Eqs. (1)-(2) is the spin temperature. This quantity is defined as $n_{1}/n_{0}\equiv 3\,\mathrm{exp}\left(-T_{\star}/T_{s}\right)$, where $n_{1}$ ($n_{0}$) denotes the number density of hydrogen atoms in the hyperfine triplet (singlet) state and $T_{\star}\equiv\hbar c/\left(k_{B}\lambda_{21}\right)\approx 68.2\,\mathrm{mK}$. In theory, the solution for $T_{s}$ has to be obtained by solving the Boltzmann equation for $n_{1}$ and $n_{0}$. Yet, we show in Appendix A that the solution for this equation is very well approximated by the well known expression for $T_{s}$ in equilibrium, even in the dark ages where $H\gg A_{10}$. This equilibrium expression reads
[…]
where $\delta_{b}$ is the baryon energy over-density and $x_{e}\equiv n_{e}/\left(n_{\mathrm{H}}+n_{\mathrm{He}}\right)$ is the free-electron fraction, with $n_{e}$, $n_{\mathrm{H}}$ and $n_{\mathrm{He}}$ the free electron, hydrogen nuclei and helium nuclei number densities, respectively. The first line in Eq. (4) represents the change in temperature in the absence of external heating sources; the first term accounts for adiabatic cooling due to the expansion of the Universe, the second term accounts for adiabatic heating due to clustering of matter and the third term accounts for adiabatic cooling/heating as a result of increasing/lowering the total number of baryons in reionization/recombination. The second line in Eq. (4) represents heating by external sources, with the first (second) term accounting for heating by CMB (X-ray) photons. The expression for Compton heating by CMB photons is well-known and can be found in many papers in the literature, including Refs. Peebles (2020); Ma and Bertschinger (1995); Seager et al. (1999, 2000); Barkana and Loeb (2005b); Naoz and Barkana (2005). In contrast, the calculation of X-ray heating, relevant only during the cosmic dawn and reionization epochs, is much more complicated and model dependent, see e.g. Ref. Mesinger et al. (2011).
II.2 Growth factor
According to the standard cosmological model, two kinds of matter controlled the evolution of the Universe at early times, before the first atoms were formed. The first kind are baryons, which include hydrogen and helium nuclei and electrons [^3] We follow standard nomenclature and cavalierly group electrons with baryons. and protons that interact strongly with photons due to their electric charge. The second kind is collisionless CDM. While the baryons were subject to radiation pressure on small scales, CDM was able to freely cluster under the influence of self-gravitational forces. As the Universe expanded, it eventually reached a low enough temperature such that neutral atoms formed for the first time. This promptly led to a dramatic increase in the photon mean free path, thus releasing the freely propagating CMB radiation. This key moment in the history of the Universe, dated at redshift $z\approx 1100$, is known as “recombination”.